# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ 6.3 Representation
We have shown so many examples with regular interior degrees that one can grow to expect that as the only possibility. The Example 6.6 is a situation where the minimum degree node $v_{0}$ has degree 6. Then in the first rooted neighborhood we have the node 1 that has two interior neighbors, $2,3$. All other nodes in that rooted neighborhood only have one interior neighbor in $R_{1}$. This is not a problem for the Decreasing Neighborhood Sequence Property though. In fact, a node relating to more nodes in an earlier container implies that it will either (1) relate to less nodes in a later container, thereby still complying with Lemma 6.5 (Decreasing Exteriors). The alternative is that a node that has a higher interior degree could have a higher overall degree. This node would still be bound though by 6.5, to prevent its parents from having their degrees double. The fact that all nodes in the first rooted neighborhood still have interior degree of at least 1 means that they are all participating in a cycle, which means that they are all contributing to the load balancing.
With this rooted neighborhood representation of the SSNC, by way of the GLOVER data structure, we can begin to visualize elements of these lemmas. For example, Lemma 6.6 (DNSP Neighborhood Size Property) can be visualized by example 6.7. This shows a series of ordered rooted neighborhoods. Each rooted neighborhood has fewer nodes. The fewer nodes imply that the rooted neighborhood circle can be drawn with a smaller radius, giving light to the decrease in size.
To further illustrate the Graph Level Order (GLOVER) data structure, imagine a social media example where users can follow one another, but no one follows back. This network is then an oriented graph. Running the SSNC in this context means looking for an account whose indirect sphere of influence is at least as large as their direct sphere of influence. Here, we are defining the indirect sphere of influence as those accounts they can reach through others. Similarly, we are defining the direct sphere of influence as those accounts they can directly connect with. Essentially, these are the same definitions as the second out-neighbor and first out-neighbor, respectively, that we have been working with. After finding the person who connects to the fewest accounts in this social media graph, we could set up neighborhoods just as before.
Now, Lemma 6.2 (Interior Load Balance) says that each person in rooted neighborhood 1 must follow another person in rooted neighborhood 1. What this is saying is that, inside this first rooted neighborhood, we will have one cycle of users where the messages are traversed about or a number of disjoint cycles where messages are spread among themselves. Consider Example 6.8, which takes the former of these approaches. These cycles in the first rooted neighborhood are what stop the user, Andi, from being an influencer.
This brings us to the next rooted neighborhood. Lemma 6.1 (Decreasing Neighborhood Sequence Property) lets us know that this rooted neighborhood will have 8 people—Kai, Lei, Mia, Nova, Owen, Pat, Qin, and Ryan. If any person in this rooted neighborhood followed less than two other people in this same rooted neighborhood, then they would need to follow more people in the next rooted neighborhood (neighborhood 3). From this we would conclude that a person in rooted neighborhood 1, whose indirect sphere of influence is this person’s sphere of influence, would have its first degree double and its second degree double, implying that its overall degree doubles.
Example 6.7.
Figure 14: Illustration of the DNSP Neighborhood Size. We have a node $x$ in the rooted neighborhood $R_{i}$ connecting to three nodes in the rooted neighborhood $R_{i+1}$. In order to prevent $x$’s degree from doubling, we need to have $|R_{i+2}|<|R_{i+1}|$.
Example 6.8.
Figure 15: This is an example of a social media setup of the SSNC. The third rooted neighborhood has 21 arcs and seven nodes, which represents an orientation of the complete graph.
@@ 7 Back Arcs
With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs. We will investigate the influence these back arcs have on our data structure and our search for a counterexample. Understanding back arcs is crucial because their presence can invalidate certain structural assumptions that underpin our search for a counterexample, including violating the principles of the partition. Their presence can create shortcuts in the graph, potentially leading to a node’s second out-neighbor overlapping with earlier rooted neighborhoods in ways that violate the distance-based ordering of GLOVER.
It is important to note that the back arcs and the exterior out-neighbors are distinct concepts. Back arcs are not a subset of exterior out-neighbors. While back arcs may connect to any previous rooted neighborhood, exterior out-neighbors are limited to connecting with the next rooted neighborhood. In addition, exterior out-neighbors do not represent all possible future rooted neighborhoods. Rather, an exterior out-neighbor $\text{ext}(x,y)$ specifically refers to nodes at a distance of two from $x$ that are also distance one from $y$. This is mainly a conceptual thing. The fact that back arcs do not point in a forward direction and exterior arcs do does not change the fact that both exterior arcs and back arcs point a parent node to a rooted neighborhood that is distance two away. Therefore, it is both reasonable and useful to treat back arcs as a subclass of exterior arcs. They should be distinguished not by their directional flow but by the fact that they cross boundaries of the rooted neighborhoods.
If we look at Example 4.2 again, remember that this example was important because of the cycle among $v_{1},v_{2},v_{3}$. We see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|\text{int}(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$. Similarly, we see that $|\text{ext}(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
One of the things we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. The next lemma formalizes that concept.
@@ 7.1 Consequences of Back Arcs
The key assumption of the GLOVER data structure is that the rooted neighborhoods are disjoint. While the rooted neighborhoods form a partition of the nodes of the graph— ensuring that each node belongs to exactly one rooted neighborhood—arcs will respect this partition. This distinction becomes very important. In theory, we do not want arcs that are able to influence the nature of the data structure. However, back arcs—edges that go from a higher-indexed neighborhood back to a lower one—can cause overlap between exterior sets. These arcs introduce connections that undermine the neat separation the rooted neighborhoods imply. As such, the validity of the lemma depends not just on the rooted neighborhood structure but on the absence or control of these back arcs.
Recall that a back arc from a node in a rooted neighborhood $R_{i}$ is an arc directed to a node in neighborhood $R_{j}$, where $j<i$. These back arcs introduce a non-empty intersection between the sets of nodes reachable from $R_{i}$ and $R_{j}$. We know that in the absence of back arcs, these rooted neighborhoods (or, more precisely, the sets of nodes relevant to our arguments) are disjoint. Therefore, we must carefully examine how the introduction of back arcs disrupts the disjoint nature and impacts our proofs. The rooted neighborhoods are members of a partition and thus are disjoint by definition.
@@
Back Arcs Disjoint Let $G$ be an oriented graph with minimum out-degree $v_{0}$ and node rooted neighborhoods $R_{i}$ and $R_{j}$, with $i>j$. If there are no back arcs from any node in $R_{i}$ to any node in $R_{j}$, then $R_{i}$ and $R_{j}$ are disjoint (i.e., $R_{i}\cap R_{j}=\emptyset$).
@@
Lemma 7.1 sets up a situation where we first establish the presence of the GLOVER data structure. We want to establish that an oriented graph without back arcs will have no intersecting rooted neighborhoods. Our proof proceeds by contradiction, supposing that a node is in two different rooted neighborhoods. The problem with this assumption is that the definition of a node belonging to a rooted neighborhood $R_{i}$ means that node is distance $i$ from the minimum out-degree node $v_{0}$. So a node $v$ being in two different rooted neighborhoods means it is at two different distances from $v_{0}$. This is not possible, though, and we conclude that they must be disjoint.
@@
Back Arcs Overlap Let $G$ be an oriented graph with minimum degree node $v_{0}$ and rooted neighborhoods $R_{i}$ and $R_{j}$, where $j<i$. If $R_{i}$ and $R_{j}$ are not disjoint, then there exists a back arc from a node in $R_{i}$ to a node in $R_{j}$.
@@
Suppose that $R_{i}$ and $R_{j}$ are not disjoint. Then there exists a node $v$ such that $v\in R_{i}\cap R_{j}$. This means that $v\in R_{i}$ and $v\in R_{j}$. Since $i<j$, and both $v\in R_{i}$ and $v\in R_{j}$, there must be a path from $v_{0}$ to $v$ of length $i$ and a path from $v_{0}$ to $v$ of length $j$. The only way this is possible is if there is a back arc from a node in $R_{i}$ to a node in $R_{j}$. Then we would extend the path from $v_{0}\rightarrow R_{i}\rightarrow v_{1}\rightarrow back\_arc\rightarrow R_{j}\rightarrow v_{2}$. Thus, the existence of non-disjoint rooted neighborhoods, when $i<j$, implies the existence of a back arc.
@@
Lemma 7.2 sets up all the possible situations in which two neighborhoods can intersect. This lemma is simply a contrapositive of lemma 7.1. There is more, though, when the back arcs are broken down as consecutive rooted neighborhoods of neighborhoods. We can see a visualization of this in 7.1. We see a situation where we have five (or more) arcs, $x\rightarrow y$, $y\rightarrow z$, $y\rightarrow\ldots$, $\ldots\rightarrow u$, $u\rightarrow w$, and $w\rightarrow z$.
@@
In 7.1, we observe that the node $w$ is distance two from $x$ and distance one from $y$, i.e., in $\text{ext}(x,y)$. The nodes preceding the back arc are $u$ and $z$, so $\text{ext}(u,z)={w}$. Without the arc $(z,w)$, the set $\text{ext}(x,y)=\emptyset$. There may be other nodes on the path from $y$ to $u$, and we can calculate those exteriors as well. The truth about any such path is that its distance to $w$ is greater than two, so $w$ would not be in the set of exterior out-neighbors. The other possibility is that a node along the path from $y$ to $u$, is in the set $\text{ext}(x,y)$. This is not possible, though,because the element that will be in $\text{ext}(x,y)$ is the first element along the path, and that cannot be the element that is in the exterior out-neighbor from the path, as it has to be distance two from this first node.

### v12

@@ 6.2 Neighborhood Density
We have shown so many examples with regular interior degrees that one can grow to expect that as the only possibility. The Example 3 is a situation where the minimum degree node $v_{0}$ has degree 6. Then in the first rooted neighborhood we have the node 1 that has two interior neighbors, $2,3$. All other nodes in that rooted neighborhood only have one interior neighbor in $R_{1}$. This is not a problem for the Decreasing Neighborhood Sequence Property though. In fact, a node relating to more nodes in an earlier container implies that it will either (1) relate to less nodes in a later container, thereby still complying with Lemma 6.6 (Decreasing Exteriors). The alternative is that a node that has a higher interior degree could have a higher overall degree. This node would still be bound though by 6.6 (Decreasing Exteriors), to prevent its parents from having their degrees double. The fact that all nodes in the first rooted neighborhood still have interior degree of at least 1 means that they are all participating in a cycle, which means that they are all contributing to the load balancing.
What we will see through the next lemma is that we can in fact use this lemma to bound the size of entire rooted neighborhoods.
Lemma 6.8.
|$|\{y\in\text{ext}(u,v)\text{ s.t. }u\in R_{i},\ v\in R_{i+1},\ y\in R_{i+2}\}|\leq d^{+}(v_{0})-i.$
Proof.
We proceed by induction on the distance $i$.
Base Case ($i=0$): The neighborhood $R_{0}=\{v_{0}\}$, and its out-neighbors form $R_{1}$, with $|R_{1}|=d^{+}(v_{0})$ by definition. Now consider $R_{2}$, which consists of nodes reachable by an arc from nodes in $R_{1}$. We assume that all nodes, including $v_{0}$ and all nodes in $R_{1}$ have the DNSP. Let $v_{0}\in R_{0}$, $v\in N^{+}(v_{0})=R_{1}$, and define $\text{ext}(v_{0},v):=N^{++}(v_{0})\cap N^{+}(v)$. From DNSP holding for $v_{0}$, we have
$|N^{++}(v_{0})|<|N^{+}(v_{0})|=d^{+}(v_{0}),$
So no matter what set we intersect $N^{++}(v_{0})$ with from $R_{1}$, the resulting intersection $N^{++}(v_{0})\cap N^{+}(v)$ will have a smaller cardinality than $d^{+}(v_{0}$.
This is the bound we need for the exterior set of $v_{0}\to v$.
$|\text{ext}(v_{0},v)|\leq d^{+}(v_{0})-1.$
Inductive Hypothesis: Assume that for some $i\geq 0$,
|$|\{y\in\text{ext}(u,v)\text{ s.t. }u\in R_{i},\ v\in R_{i+1},\ y\in R_{i+2}\}|\leq d^{+}(v_{0})-i.$
Inductive Step: Consider $u\in R_{i+1}$, and $u\to v\in G$, where $v\in R_{i+2}$, and we will examine $\text{ext}(u,v):=N^{++}(u)\cap N^{+}(v)$.
From Lemma 6.6 (Decreasing Exteriors Lemma), we have:
$|\text{ext}(u,v)|<|\text{ext}(w,u)|,$
for any parent $w\in R_{i}$, $u\in R_{i+1}$.
By the inductive hypothesis,
$|\text{ext}(w,u)|\leq d^{+}(v_{0})-i.$
We can combine these inequalities to get,
$|\text{ext}(u,v)|<d^{+}(v_{0})-i\to|\text{ext}(u,v)|\leq d^{+}(v_{0})-(i+1).$
What this has established through the DNSP, and contingent upon the proof of the generalization of load balancing, is that the rooted neighborhoods are decreasing in size as we move further away from the minimum out-degree node $v_{0}$. We spoke of the DNSP being a bottleneck, and here is another example where, across the entire graph, we are actually seeing it restrict the sizes of rooted neighborhoods. DNSP is a local bottleneck that induces a global contraction of the neighborhood layers in the Graph Level Order.
We began with the assumption that the Decreasing Neighborhood Sequence Property (DNSP) established that for a node $u$, the size of its first out-neighborhood is strictly larger than its second out-neighborhood $|N^{+}(u)|>|N^{++}(u)|$. We have now witnessed the emergence of this principle at the level of entire rooted neighborhoods within the Graph Level Order structure. Lemma 6.8 (DNSP Impact on Rooted Neighborhood Size) shows that consecutive neighborhoods must strictly decrease in size $(|R_{i+1}|<|R_{i}|$ for $i\geq 1)$. This demonstrates DNSP rising from a local node-level constraint to a global feature of the graph’s partitioning. This highlights the penetrating and fundamental nature of the DNSP. It is felt at every node of the graph, impacting not only the individual nodes but the hierarchies above them as well. Because the Graph Level Order is organized in this way, it merges the power of the DNSP to truly organize these levels. This rooted neighborhood-level constraint of the DNSP sets the stage for the next level of understanding between the implications for the existence of back arcs. These back arcs hold weight in whether the Graph Level Order is just another data structure or if it can help solve the SSNC.
We have shown $|R_{i+1}|<|R_{i}|$, which means $|R_{i+1}|\leq|R_{i}|-1$. Now that we know that the rooted neighborhoods decrease in size, we can quantify the bounds on these sizes based on the distances from the minimum out-degree node. This will help with factors like back arcs and the ability to generalize the above lemma.
Lemma 6.9.
$|R_{0}|=1,$
$|R_{1}|=\delta,$
$|R_{i}|\leq\delta-(i-1)$
Proof.
We proceed by induction on the distance $i$ of rooted neighborhoods $R_{i}$.
Base cases:
  • $R_{0}=\{v_{0}\}$, so $|R_{0}|=1$ by definition.
  • $R_{1}=N^{+}(v_{0})$, so $|R_{1}|=\delta$, again by definition.
The DNSP says $|N^{+}(v_{0})|>|N^{++}(v_{0})|$. All of the out-neighbors of $v_{0}$ are in the rooted neighborhood $R_{1}$ and every second neighbor of $v_{0}$ will be in the rooted neighborhood $R_{2}$. By Lemma 6.2 (Interior Load Balancing), these children $u\in R_{1}$ help prevent $v_{0}$ from becoming a degree doubling node by forming cycle(s) in $R_{1}$. This bounds the number of external neighbors in $R_{2}$ and thus the size of $|R_{2}|$.
Inductive hypothesis: Assume that for some $k<i$, we have:
$|R_{k}|\leq\delta-(k-1)$
Inductive step: We consider the neighborhood $R_{k+1}$. We apply Lemma 6.8 (DNSP Impact on Neighborhood Size), which implies:
$|R_{k+1}|<|R_{k}|\leq\delta-(k-1)\to|R_{k+1}|\leq\delta-((k+1))-1)=\delta-k.$
We conclude that the inductive structure holds for all $i$, giving:
$|R_{i}|\leq\delta-(i-1),$
What we saw from 6.8 (DNSP Impact on Rooted Neighborhood Size) is that the rooted neighborhoods decreased in size. This decrease must be at least by one at each neighborhood. What 6.9 (Formula for Rooted Neighborhood Size) does is combine this into a worse case calculation size for each rooted neighborhood. In addition, this allows us to calculate the worst case sizes for interior and exterior neighborhood sizes. This gives us a way to evaluate a node’s second neighborhood against these rooted neighborhood sizes, as a comparison for degree-doubling nodes.
@@
What we also saw in this section, though, is that as this load is passed from parent to child the size of these neighborhoods is reduced. This leads to a sequence of rooted neighborhoods that are getting smaller and smaller, but these smaller neighborhoods are expected to hold more nodes inside them. Obviously there is a collision that is inevitable. We need to ensure that other concepts like back arcs do not get in the way. The Decreasing Neighborhood Sequence Property (DNSP) says that we should continuously have these rooted neighborhoods decrease forever. Any time that they stop decreasing, we will arrive at a node, or set of nodes, whose degree doubles. This would be a violation of the DNSP assumption and a contradiction.
@@ 7 Back Arcs
With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs. We will investigate the influence these back arcs have on our data structure and our search for a counterexample. Understanding back arcs is crucial because their presence can invalidate certain structural assumptions that underpin our search for a counterexample, including violating the principles of the partition. Their presence can create shortcuts in the graph, potentially leading to a node’s second out-neighbor overlapping with earlier rooted neighborhoods in ways that violate the distance-based ordering of the Graph Level Order.
It is important to note that the back arcs and the exterior out-neighbors are distinct concepts. Back arcs are not a subset of exterior out-neighbors. While back arcs may connect to any previous rooted neighborhood, exterior out-neighbors are limited to connecting with the next rooted neighborhood. In addition, exterior out-neighbors do not represent all possible future rooted neighborhoods. Rather, an exterior out-neighbor $\text{ext}(x,y)$ specifically refers to nodes at a distance of two from $x$ that are also distance one from $y$. However, this is mainly a conceptual thing. The fact that back arcs do not point in a forward direction and exterior arcs do, does not change the fact that both exterior arcs and back arcs point a parent node to a node in a different rooted neighborhood. Therefore, it is both reasonable and useful to treat back arcs as a subclass of exterior arcs. They should be distinguished not by their directional flow but by the fact that they cross boundaries of the rooted neighborhoods.
If we look at Example 4.2 again, remember that this example was important because of the cycle among $v_{1},v_{2},v_{3}$. We see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|\text{int}(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $v_{1}\to v_{2}\in G$. Similarly, we see that $|\text{ext}(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $v_{1}\to v_{4}\in G$ and $v_{1}\to v_{5}\in G$ and $v_{0}\to v_{4}\in G$ and $v_{0}\to v_{5}\in G$.
Something we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. The next lemma formalizes that concept.
@@ 7.1 Consequences of Back Arcs
The key assumption of the Graph Level Order is that the rooted neighborhoods are disjoint. While the rooted neighborhoods form a partition of the nodes of the graph— ensuring that each node belongs to exactly one rooted neighborhood—arcs will respect this partition. This distinction becomes very important. In theory, we do not want arcs that are able to influence the nature of the data structure. However, back arcs—edges that go from a higher-indexed neighborhood back to a lower one—can cause overlap between exterior sets. These arcs introduce connections that undermine the neat separation the rooted neighborhoods imply. As such, the validity of the lemma depends not just on the rooted neighborhood structure but on the absence or control of these back arcs.
Recall that a back arc from a node in a rooted neighborhood $R_{i}$ is an arc directed to a node in neighborhood $R_{j}$, where $j<i$. These back arcs introduce a non-empty intersection between the sets of nodes reachable from $R_{i}$ and $R_{j}$. In the absence of back arcs, these rooted neighborhoods (or, more precisely, the sets of nodes relevant to our arguments) are disjoint. Therefore, we must carefully examine how the introduction of back arcs disrupts the disjoint nature and impacts our proofs. The rooted neighborhoods are members of a partition and thus are disjoint by definition.
@@
Back Arcs Necessary Let $G$ be an oriented graph with minimum out-degree $v_{0}$ and node rooted neighborhoods $R_{i}$ and $R_{j}$, with $i>j$. If there are no back arcs from any node in $R_{i}$ to any node in $R_{j}$, then $R_{i}$ and $R_{j}$ are disjoint (i.e., $R_{i}\cap R_{j}=\emptyset$).
@@
Lemma 7.1 sets up a situation where we first establish the presence of the Graph Level Order. We want to establish that an oriented graph without back arcs will have no intersecting rooted neighborhoods. Our proof proceeds by contradiction, supposing that a node is in two different rooted neighborhoods. The problem with this assumption is that the definition of a node belonging to a rooted neighborhood $R_{i}$ means that node is distance $i$ from the minimum out-degree node $v_{0}$. So a node $v$ being in two different rooted neighborhoods means it is at two different distances from $v_{0}$. This is not possible, though, and we conclude that they must be disjoint.
@@
Back Arcs Sufficient Let $G$ be an oriented graph with minimum degree node $v_{0}$ and rooted neighborhoods $R_{i}$ and $R_{j}$, where $j<i$. If $R_{i}$ and $R_{j}$ are not disjoint, then there exists a back arc from a node in $R_{i}$ to a node in $R_{j}$.
@@
Suppose that $R_{i}$ and $R_{j}$ are not disjoint. Then there exists a node $v$ such that $v\in R_{i}\cap R_{j}$. This means that $v\in R_{i}$ and $v\in R_{j}$. Since $i<j$, and both $v\in R_{i}$ and $v\in R_{j}$, there must be a path from $v_{0}$ to $v$ of length $i$ and a path from $v_{0}$ to $v$ of length $j$. The only way this is possible is if there is a back arc from a node in $R_{i}$ to a node in $R_{j}$. Then we would extend the path from $v_{0}\to R_{i}\to v_{1}\to back\_arc\to R_{j}\to v_{2}$. Thus, the existence of non-disjoint rooted neighborhoods, when $i<j$, implies the existence of a back arc.
@@
Lemma 7.2 (Back Arcs Sufficient) sets up the necessary conditions in which two neighborhoods can intersect. This lemma is simply a contrapositive of Lemma 7.1 (Back Arcs Required). We can see a visualization of this in Example 7.1. We see a situation where we have five (or more) arcs, $y\to w$, $x\to y\to\ldots\to u\to z\to w$.
@@
In Example 7.1, we observe that the node $w$ is distance two from $x$ and distance one from $y$, i.e., in $\text{ext}(x,y)$. The nodes preceding the back arc are $u$ and $z$, so $\text{ext}(u,z)=\{w\}$. Without the arc $z\to w$, the set $\text{ext}(x,y)=\emptyset$. There may be other nodes on the path from $y$ to $u$, and we can calculate those exteriors as well. The truth about any such path is that its distance to $w$ is greater than two, so $w$ would not be in the set of exterior out-neighbors. The other possibility is that a node along the path from $y$ to $u$, is in the set $\text{ext}(x,y)$. This is not possible, though,because the element that will be in $\text{ext}(x,y)$ is the first element along the path, and that cannot be the element that is in the exterior out-neighbor from the path, as it has to be distance two from this first node.